OpenStudy (anonymous):

Help Me :D -y/2 (<=) 3 A) y (<=) 6 B) y (<=) -6 C) y (>=) -6 D) y (>=) 6

OpenStudy (anonymous):

When you multiply by a negative (how you're going to solve it) you must turn around the inequality sign

Nnesha (nnesha):

first you have to multiply both side by 2 and then divide both side by -1 and remember when you multiply or divide by negative number you must have to flip the sing

OpenStudy (anonymous):

\[\Large \frac{-y}{2}\le3\]

Nnesha (nnesha):

sign**

Nnesha (nnesha):

let us know what you get ??

OpenStudy (anonymous):

yea thats it @doulikepiecauseidont

OpenStudy (anonymous):

\[\Large 2 \times \left( \frac{-y}{2} \right) \le 3 \times 2 \]

OpenStudy (anonymous):

\[\Large (-1) \times -y \le6 \times (-1)\]

OpenStudy (anonymous):

\[\Large y \ge -6\]

OpenStudy (anonymous):

so C right ?